OpenStudy (anonymous):

Last one Consider the function f(x) = 8/x^{3}-4/x^{5}. Let F(x) be the antiderivative of f(x) with F(1) = 0. Then F(5) equals

OpenStudy (anonymous):

f(x)=\[\frac{ 8 }{ x^3 }-\frac{ 4 }{ x^5 }\]

OpenStudy (anonymous):

\[8x ^{-3}\] \[4x ^{-5}\]

OpenStudy (anonymous):

\[4x ^{-2}\] or \[\frac{ 4 }{ x^2 }\] and \[x ^{-4}\] or\[\frac{ 1 }{ x^2 }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ x^4 }\]

OpenStudy (anonymous):

\[F(x)=\frac{ 1 }{ x^4 }-\frac{ 4 }{ x^2 }\]